# LEASE: Offline Preference-based Reinforcement Learning with High Sample Efficiency

### v1

@@ Abstract
Offline preference-based reinforcement learning (PbRL) provides an effective way to overcome the challenges of designing reward and the high costs of online interaction. However, since labeling preference needs real-time human feedback, acquiring sufficient preference labels is challenging. To solve this, this paper proposes a offLine prEference-bAsed RL with high Sample Efficiency (LEASE) algorithm, where a learned transition model is leveraged to generate unlabeled preference data. Considering the pretrained reward model may generate incorrect labels for unlabeled data, we design an uncertainty-aware mechanism to ensure the performance of reward model, where only high confidence and low variance data are selected. Moreover, we provide the generalization bound of reward model to analyze the factors influencing reward accuracy, and demonstrate that the policy learned by LEASE has theoretical improvement guarantee. The developed theory is based on state-action pair, which can be easily combined with other offline algorithms. The experimental results show that LEASE can achieve comparable performance to baseline under fewer preference data without online interaction.
Index Terms:
Preference-based reinforcement learning; Sample efficiency; Reward learning; Theoretical guarantee
@@ I Introduction
Reinforcement learning (RL) has been widely applied in robot control [1, 10] and healthcare [12, 11] fields. However, in the real-world scenarios, RL faces two serious challenges: 1) Since RL learns the optimal policy through trial and error, it is dangerous for agent to interact directly with environment, especially human-in-loop control [40]; 2) Due to the complexity of human cognitive and control system, it is difficult to design a quantitative reward function to accurately reflect the human intention or preference [9]. Many researches have been explored to solve the above two challenges.
Offline RL, learning policy from previous collected dataset without interaction with environment, has become an effective way for challenge 1 [4, 8]. However, due to limited data coverage, offline RL suffers from the distribution shift between the learned policy and the behavior policy [28], which leads to poor agent performance. To address this problem, offline RL either directly constrain the policy closed to the behavior policy [2] or penalize value function for out-of-distribution (OOD) data [4, 8]. The above methods improve the performance of agent through incorporating conservatism into policy learning.
Preference-based RL (PbRL), also known as RL from human feedback (RLHF), which learns the reward function from human preference data without the demand for tedious hand-engineered reward design [14, 43], has been developed recently for challenge 2. However, compared with data $(s,a,s^{\prime})$, the collecting cost of preference data is higher since it demands human real-time feedback to label preference, which often brings high sample complexity (low sample efficiency) [13, 3]. Some data augmentation techniques, such as generating unlabeled data through interaction with simulator [3] or temporal cropping [35], have been proposed to solve problem of the limitation of preference data. However, the above methods need online interaction with environment (simulator) and still demand large human feedback. Thus, in PbRL, the sample efficiency is still of utmost importance.
In offline PbRL setting, existing methods [34, 17, 33] typically involve two steps: reward learning and policy learning. They learns the policy based on offline static dataset without reward and preference dataset, and introduce transformer [17] or diffusion model [33] to improve the performance of reward model and agent performance. However, the theoretical guarantees of reward generalization and policy improvement are not provided. Therefore, this paper aims to design a high sample efficiency offline PbRL algorithm that can achieve comparable performance with baseline from fewer preference data without interaction with environment, and develop the related theory for reward generalization bound and policy improvement guarantee. The key questions can be summarized as follow:
@@
    In algorithm, the collected preference data fails to cover all state-action region and the collecting cost is too high. How to use the limited preference data to learn accurate reward model and guarantee agent performance?
@@
    In theory, what factors affect the reward generalization and what is the detailed relationship between offline RL algorithm itself, the reward gap and the improvement of policy performance?
For question 1, motivated by model-based offline RL, which learns the environment model to broaden data coverage [37, 36], we also train the transition model to achieve data augmentation (improve sample efficiency). Specifically, the agent generates two different trajectories through interaction with the learned transition model, and then the pretrained reward model generates pseudo label for these two trajectories. However, the preference model may generate incorrect pseudo label, which may bring unstable training and low performance.
For question 2, there are very fewer algorithms for offline PbRL theory. The most relevant work is [18], where the theory of policy improvement guarantee is established. However, the generalization bound for reward model is not considered, and this theory is based on the whole trajectory rather than state-action pair in offline algorithms, which is not conducive to the theoretical analysis of specific offline PbRL algorithms since many offline RL are based on state-action pair.
Figure 1: The learning framework of LEASE. The offline dataset without reward is used to train transition model and limited labeled preference dataset is used to pretrain reward model. Then the generated unlabeled preference data is screened through uncertainty-aware selecting mechanism. The reward model is updated based on labeled and generated dataset. Finally, the agent learns the policy based on offline dataset and learned reward model.
To solve the above problems, this paper proposes a novel offLine prEference-bAsed RL with high Sample Efficiency algorithm (LEASE), where a selecting mechanism is designed to guarantee the quantity of generated dataset. This selecting mechanism is based on the principles of high confidence and low uncertainty, and can effectively reduce the error of pseudo-labels. Moreover, we develop a new generalization bound for reward model, and provide the theory of policy improvement guarantee for offline PbRL based on state-action pair. Fig. 1 shows the learning framework of LEASE. The theoretical and experimental results show that LEASE has policy improvement guarantee and can achieve comparable performance to baseline under fewer preference data on D4RL benchmark. The contributions are given below:
@@
    An uncertainty-aware mechanism is designed for screening the generated preference data so as to guarantee the stability of reward training and improve the accuracy of reward prediction.
@@
    The generalization bound of reward model and the theory of policy improvement based on state-action pair are developed to analyze factors that influence reward and policy performance.
@@
The framework of this paper is as follows: Section II introduces the related works about offline PbRL. Section III provides the basic notations for algorithm. Section IV illustrates the two main problems that this paper aims to solve. Section V introduces the details of designed selecting mechanism and derives the generalization bound for reward model. Section VI illustrates the implementation details of the proposed method and provides theory for offline PbRL. Section VII conducts experiments to validate the effectiveness of LEASE and analyze the effects of each part for agent performance. Section VIII further discusses the offline PbRL theory and algorithm. Section IX summarizes the entire work.
II Related Work
@@
This section presents a brief overview for the algoriths and related theories of offline PbRL, and the differences and advantages between our work and the previous related works.
@@
The algorithm for offline PbRL. Offline PbRL eliminates the demand for interaction with environment and handcraft designed reward. OPAL [34] is the first algorithm that combining offline RL and PbRL. PT [17] utilized transformer-based architecture to design preference model capable of generating non-Markovian rewards. OPPO [30] directly optimized the policy in a high-level embedding space without learning a separate reward function. However, the above methods demands a large amount of preference dataset. FTB [33] trains a diffusion model to achieve data augmentation, but this consumers larger training cost. Instead, our method LEASE trains simple transition model to generate unlabeled data and designed selecting mechanism to ensure the quality of the generated data, which can achieve superior performance under fewer dataset and less time.
@@
The theory for offline PbRL. There are few algorithms that provide theoretical guarantees for offline PbRL, including the generalization bound of reward model and the guarantee of policy improvement. Zhu et al.[27] studied offline PbRL, but the analysis are restricted to linear model. Zhan et al.[18] extended to general function approximation, but the generalization analysis of reward model is not provided and the theory is based on trajectory. LEASE gives the theoretical analysis for reward model, and provides the theoretical guarantee for policy based on state-action pair. The generalization bound of reward model has generality and can be applicable to methods that train reward model using pseudo-labels, and the theory of policy improvement guarantee of offline PbRL can be easily combined with other offline RL.
@@
Data selection techniques. Data screening mechanism essentially belongs to the field of semi-supervised learning. The goal of semi-supervised learning is using unlabeled data to improve model performance when labeled data is limited. Data selection techniques is used to filter the data with clean labels from a noisy dataset. Han et al.[7] selected unlabeled data with small losses from one network. Li et al.[5] used a two-component GMM to separate the dataset into a clean set and a noisy set. Rizve et al.[31] modeled the prediction uncertainty of unlabeled data to screen data. Xiao et al.[6] selected data with high confidence as clean data. Motivated by [31] and [6], we trained ensemble reward model to select data with high confidence and low variance to guarantee the quality of unlabeled data.
In PbRL methods, Surf [3] is the most similar to our method LEASE that uses data augmentation technique. The differences between them mainly includes the below three aspects: 1) Surf belongs to online RL, but LEASE belongs to offline RL. Surf generates data through interaction with environment (simulator) while LEASE generates data through the learned transition model. Note that, in most human-in-loop-control scenarios, the cost of online interaction is large and the realistic simulator is difficult to create; 2) Surf only uses confidence for label filtering, whereas LEASE employs both confidence and uncertainty principles, which effectively reduce pseudo-label error. The results in Table III have validated this; 3) LEASE provides the general theoretical framework for offline PbRL, but Surf don’t provide theoretical analysis.
III Preliminaries
In this section, we provides the basic notations for reinforcement learning algorithm, including offline RL, preference-based RL and model-based RL.
@@ III Preliminaries
Offline reinforcement learning. The framework of RL is based on the Markov Decision Process (MDP) that is described by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},R,T,\rho,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta({\mathcal{S}})$ is the transition dynamics, $\rho$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor [40]. The term $\Delta({\Omega})$ denotes the set of probability distribution over space $\Omega$. The goal of RL is to optimize the policy $\pi$ that maximizes the expected discounted return $J(\pi,R):=\mathbb{E}_{(s,a)\sim d_{T}^{\pi}(s,a)}[R(s,a)]/(1-\gamma)$, where $d_{T}^{\pi}(s,a):=d_{T}^{\pi}(s)\pi(a|s)$ is the state-action marginal distribution under the learned policy $\pi$. The discounted state marginal distribution $d_{T}^{\pi}(s)$ is denoted as $(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathcal{P}(s_{t}=s|\pi)$, where $\mathcal{P}(s_{t}=s|\pi)$ is the probability of reaching state $s$ at time $t$ by rolling out $\pi$. The policy $\pi$ can be derived from $Q$-learning [41], which learns the $Q$-function that satisfies Bellman operator: $\mathcal{T}Q(s,a):=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim T_{\mathcal{M}}(s^{\prime}\mid s,a)}[\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})]$.
@@
Preference-based reinforcement learning. Different from standard RL setting, the reward function is not available in PbRL. Instead, PbRL learns the reward function $\widehat{R}$ from preferences between pairs of trajectory segments to align the human intention [15], where a trajectory segment $\sigma$ of length $L$ is defined as a sequence of states and actions $\{s_{k},a_{k},...,s_{k+L-1},a_{k+L-1}\}\in(\mathcal{S}\times\mathcal{A})^{L}$. Given a pair of segments $(\sigma_{0},\sigma_{1})$, human choose which segment is preferred, i.e., $y\in\{0,1\}$ [14]. The preference label $y=1$ indicates $\sigma_{1}\succ\sigma_{0}$ and $y=0$ indicates $\sigma_{0}\succ\sigma_{1}$, where $\sigma_{i}\succ\sigma_{j}$ denotes that the segment $i$ is preferable to the segment $j$.
@@
Model-based reinforcement learning. Model-based RL learns the dynamics model to improve sample efficiency [38, 36, 39]. They used the offline dataset $\mathcal{D}_{\text{offline }}$ to estimate transition model $\widehat{T}(s^{\prime}|s,a)$. The transition model $\widehat{T}_{\varphi}$ parameterized by $\varphi$ is typically trained via maximum likelihood estimation (MLE):
@@
In practical implantation, the transition model is approximated by Gaussian distribution and the MLE method is employed to train $N_{T}$ ensemble transition models $\{T^{i}_{\varphi}=\mathcal{N}(\mu^{i}_{\varphi},\sigma^{i}_{\varphi})\}_{i=1}^{N_{T}}$ [37]. The samples are generated through $H$-step rollouts. Here, we use the trained transition to generate more trajectory to achieve data augmentation.
IV Problem Formulation
The cost of collecting preference data is high since it needs real-time human feedback. Therefore, this paper aims to improve sample efficiency to learn accurate reward model from limited preference data and guarantee agent performance. For PbRL, the ideal form is that the learned reward function $\widehat{R}$ from collected preference data can be consistent with true reward $R^{*}$. Here, we define the function class as =$\mathcal{R}=\{R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$. Then we assume the reward class $\mathcal{R}$ is realizable.
@@ IV Problem Formulation
Based on the above assumption, the optimal reward function can be obtained if the preference data can cover all data space. However, preference data tend to be limited, even less than hundred in real-world scenarios, such as rehabilitation robot field. Therefore, different from previous PbRL [3, 35, 33], this paper learns reward function from the limited dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$. We train the transition model through Eq. (4) to generate more unlabeled preference data $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$. The pseudo labels $\widehat{y}$ for unlabeled dataset $\mathcal{D}_{u}$ can be obtained through reward model trained on $\mathcal{D}_{l}$ [3], that is
@@
where $\mathbbm{1}(\cdot)$ is indicator function.The reward function is updated through collected labeled dataset $\mathcal{D}_{l}$ and generated unlabeled dataset $\mathcal{D}_{u}$. Then, the reward model can be optimized through minimizing
@@
$+$ $\mathbb{E}_{(\sigma_{0}^{u},\sigma_{1}^{u})\sim\mathcal{D}_{u}}\big{[}L\left((\sigma_{0}^{u},\sigma_{1}^{u}),\widehat{y}\right)\big{]},$
where $L((\sigma_{0},\sigma_{1}),y)=-(1-y)\log P(\sigma_{0}\succ\sigma_{1};{\psi})-y\log P(\sigma_{1}\succ\sigma_{0};{\psi})$. However, the pretrained reward model may generate incorrect pseudo-labels for unlabeled dataset, leading to noisy training and poor generalization [31]. Therefore, one key question is how to design a data selecting mechanism to improve prediction accuracy and guarantee training stability.
The another key aspect is the theory for offline PbRL. There are very few algorithms specifically designed for offline PbRL with strong guarantee, including generalization bound for reward model and safe improvement guarantee for policy learning. Therefore, another key question is to develop a systematic theory for offline PbRL, including generalization bound and improvement guarantee.
V Reward Learning
The reward learning involves two stages: pretraining based on collected labeled data $\mathcal{D}_{l}$ and updating based on $\mathcal{D}_{l}$ and unlabeled data $\mathcal{D}_{u}$ during policy learning, which is essentially semi-supervised learning [26]. This section focuses on designing a data selecting function $f(\sigma_{0}^{u},\sigma_{1}^{u})$ to ensure the quality of generated preference data, and explaining the factors that influence the generalization ability of reward model.
V-A Uncertainty-aware Pseudo-label Selection
Motivated by the previous pseudo-labeling work [31], we select data from unlabeled dataset $\mathcal{D}_{u}$ according two principles: confidence and uncertainty. The data with high confidence and low uncertainty can be chosen for reward training. High confidence refers that pre-trained reward model discriminates the preference of two trajectories into $\widehat{y}$ with high probability $p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})$. Low uncertainty refers that the $N_{R}$ reward models (model ensembles) predicts with small variance $\tau(\sigma_{0}^{u},\sigma_{1}^{u},N_{R})$. The probability confidence $p(\sigma_{0}^{u},\sigma_{1}^{u},\widehat{y})$ and uncertainty variance $\tau(\sigma_{0}^{u},\sigma_{1}^{u},N_{R})$ can be denoted as

### v2

@@ Abstract
Offline preference-based reinforcement learning (PbRL) offers an effective approach to addressing the challenges of designing rewards and mitigating the high costs associated with online interaction. However, since labeling preference needs real-time human feedback, acquiring sufficient preference labels is challenging. To solve this, this paper proposes an offLine prEference-bAsed RL with high Sample Efficiency (LEASE) algorithm, where a learned transition model is leveraged to generate unlabeled preference data. Considering the pretrained reward model may generate incorrect labels for unlabeled data, we design an uncertainty-aware mechanism to ensure the performance of the reward model, where only high-confidence and low-variance data are selected. Moreover, the generalization bound of the reward model is provided to analyze the factors influencing reward accuracy, and the policy learned by LEASE has a theoretical improvement guarantee. The above developed theory is based on a state-action pair, which can be easily combined with other offline algorithms. The experimental results show that LEASE can achieve comparable performance to the baseline under fewer preference data without online interaction.
@@ I Introduction
Reinforcement learning (RL) has been widely applied in robot control [1, 10] and healthcare [12, 11] fields. However, in the real-world scenarios, RL faces two serious challenges: 1) Since RL learns the optimal policy through trial and error, it is dangerous for agent to interact directly with environment, especially human-in-loop control [39]; 2) Due to the complexity of human cognitive and control system, it is difficult to design a quantitative reward function to accurately reflect the human intention or preference [9]. Much research has been conducted to solve the above two challenges.
Offline RL, learning policy from previously collected dataset without interaction with the environment, has become an effective way for challenge 1 [4, 8]. However, due to limited data coverage, offline RL suffers from the distribution shift between the learned policy and the behavior policy [28, 48], which leads to poor agent performance. To address this problem, offline RL either directly constrains the policy close to the behavior policy [2, 47] or penalizes the value function for out-of-distribution (OOD) data [4, 8, 46]. The above methods improve the performance of the agent by incorporating conservatism into policy learning.
Preference-based RL (PbRL), also known as RL from human feedback (RLHF), which learns the reward function from human preference data without the demand for tedious hand-engineered reward design [14, 42], has been developed recently for challenge 2. However, compared with data $(s,a,s^{\prime})$, the collecting cost of preference data is higher since it demands human real-time feedback to label preference, which often brings high sample complexity (low sample efficiency) [13, 3]. Some data augmentation techniques, such as generating unlabeled data through interaction with a simulator [3] or temporal cropping [35], have been proposed to solve the problem of the limitation of preference data. However, the above methods need online interaction with the environment (simulator) and still demand large human feedback. Thus, in PbRL, the sample efficiency is still of utmost importance.
In offline PbRL settings, existing methods [34, 17, 33] typically involve two steps: reward learning and policy learning. They learns the policy based on an offline static dataset without a reward and preference dataset, and introduce transformer [17] or diffusion model [33] to improve the performance of the reward model and agent performance. However, the theoretical guarantees of reward generalization and policy improvement are not provided. Therefore, this paper aims to design a high sample efficiency offline PbRL algorithm that can achieve comparable performance with the baseline from fewer preference data without interaction with the environment, and develop the related theory for reward generalization bound and policy improvement guarantee. The key questions can be summarized as follows:
@@
    In the algorithm, the collected preference data fails to cover all state-action regions, and the collection cost is too high. How to use the limited preference data to learn an accurate reward model and guarantee agent performance?
@@
    In theory, what factors affect the reward generalization, and what is the detailed relationship between the offline RL algorithm itself, the reward gap, and the improvement of policy performance?
For question 1, motivated by model-based offline RL, which learns the environment model to broaden data coverage [37, 36], we also train the transition model to achieve data augmentation (improve sample efficiency). Specifically, the agent generates two different trajectories through interaction with the learned transition model, and then the pretrained reward model generates pseudo-labels for these two trajectories. However, the preference model may generate incorrect pseudo-labels, which may lead to unstable training and low performance. For question 2, there are very few algorithms for offline PbRL theory. The most relevant work is [18], where the theory of policy improvement guarantee is established. However, the generalization bound for the reward model is not considered, and this theory is based on the whole trajectory rather than the state-action pair in offline algorithms. This is not conducive to the theoretical analysis of specific offline PbRL algorithms, since many offline RL algorithms are based on the state-action pair.
Figure 1: The learning framework of LEASE. The offline dataset without reward is used to train the transition model, and a limited labeled preference dataset is used to pretrain the reward model. Then, the generated unlabeled preference data is screened through the uncertainty-aware selection mechanism. The reward model is updated based on the labeled and generated dataset. Finally, the agent learns the policy based on an offline dataset and a learned reward model.
To solve the above problems, this paper proposes a novel offLine prEference-bAsed RL with a high Sample Efficiency algorithm (LEASE), where a selecting mechanism is designed to guarantee the quantity of the generated dataset. This selection mechanism is based on the principles of high confidence and low uncertainty, and can effectively reduce the error of pseudo-labels. Moreover, we develop a new generalization bound for the reward model and provide a theory of policy improvement guarantee for offline PbRL based on the state-action pair. Fig. 1 shows the learning framework of LEASE. The theoretical and experimental results show that LEASE has a policy improvement guarantee and can achieve comparable performance to the baseline under fewer preference data on the D4RL benchmark. The contributions are given below:
@@
    An uncertainty-aware mechanism is designed for screening the generated preference data to guarantee the stability of reward training and improve the accuracy of reward prediction.
@@
    The generalization bound of the reward model and the theory of policy improvement based on the state-action pair are developed to analyze factors that influence reward and policy performance.
@@
The framework of this paper is as follows: Section II introduces the related works about offline PbRL. Section III provides the basic notations for the algorithm and illustrates the two main problems that this paper aims to solve. Section IV introduces the details of the designed selection mechanism and derives the generalization bound for the reward model. Section V illustrates the implementation details of the proposed method and provides theory for offline PbRL. Section VI conducts experiments to validate the effectiveness of LEASE and analyze the effects of each part on agent performance. Section VII further discusses the offline PbRL theory and algorithm. Section VIII summarizes the entire work.
II Related Works
@@
This section presents a brief overview of the algorithms and related theories of offline PbRL, and discusses the differences and advantages between our work and the previous related works.
@@
The algorithm for offline PbRL. Offline PbRL eliminates the demand for interaction with the environment and handcrafted-designed reward. OPAL [34] is the first algorithm that combines offline RL and PbRL. PT [17] utilized a transformer-based architecture to design a preference model capable of generating non-Markovian rewards. OPPO [30] directly optimized the policy in a high-level embedding space without learning a separate reward function. However, the above methods demand a large amount of preference data. FTB [33] trains a diffusion model to achieve data augmentation, but this consumes a larger training cost. Instead, our method LEASE trains a simple transition model to generate unlabeled data and designs a selection mechanism to ensure the quality of the generated data, which can achieve superior performance under fewer datasets and less time.
@@
The theory for offline PbRL. There are a few algorithms that provide theoretical guarantees for offline PbRL, including the generalization bound of the reward model and the guarantee of policy improvement. Zhu et al.[27] studied offline PbRL, but the analysis is restricted to the linear model. Zhan et al.[18] extended to general function approximation, but the generalization analysis of the reward model is not provided, and the theory is based on the trajectory. LEASE gives the theoretical analysis for the reward model, and provides the theoretical guarantee for the policy based on the state-action pair. The generalization bound of the reward model has generality. It can be applied to methods that train a reward model using pseudo-labels, and the theory of policy improvement guarantees in offline PbRL can be easily combined with other offline RL approaches.
@@
Data selection techniques. The data screening mechanism essentially belongs to the field of semi-supervised learning. The goal of semi-supervised learning is to use unlabeled data to improve model performance when labeled data is limited. Data selection techniques are used to filter the data with clean labels from a noisy dataset. Han et al.[7] selected unlabeled data with small losses from one network. Li et al.[5] used a two-component GMM to separate the dataset into a clean set and a noisy set. Rizve et al.[31] modeled the prediction uncertainty of unlabeled data to screen data. Xiao et al.[6] selected data with high confidence as clean data. Motivated by [31] and [6], we trained an ensemble reward model to select data with high confidence and low variance to guarantee the quality of unlabeled data.
III Preliminaries and Problem Formulation
In this section, we provide the basic notations for reinforcement learning algorithms, including offline RL, preference-based RL, and model-based RL. Then, we introduce the two main problems that this paper aims to solve from both algorithmic and theoretical perspectives.
III-A Preliminaries for RL
@@
Offline reinforcement learning. The framework of RL is based on the Markov Decision Process (MDP) that is described by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},R,T,\rho,\gamma)$, where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function, $T:\mathcal{S}\times\mathcal{A}\rightarrow\Delta({\mathcal{S}})$ is the transition dynamics, $\rho$ is the initial state distribution, and $\gamma\in(0,1)$ is the discount factor [39]. The term $\Delta({\Omega})$ denotes the set of probability distributions over space $\Omega$. The goal of RL is to optimize the policy $\pi$ that maximizes the expected discounted return $J(\pi,R):=\mathbb{E}_{(s,a)\sim d_{T}^{\pi}(s,a)}[R(s,a)]/(1-\gamma)$, where $d_{T}^{\pi}(s,a):=d_{T}^{\pi}(s)\pi(a|s)$ is the state-action marginal distribution under the learned policy $\pi$. The discounted state marginal distribution $d_{T}^{\pi}(s)$ is denoted as $(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}\mathcal{P}(s_{t}=s|\pi)$, where $\mathcal{P}(s_{t}=s|\pi)$ is the probability of reaching state $s$ at time $t$ by rolling out $\pi$. The policy $\pi$ can be derived from $Q$-learning [40], which learns the $Q$-function that satisfies Bellman operator: $\mathcal{T}Q(s,a):=R(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim T_{\mathcal{M}}(s^{\prime}\mid s,a)}[\max_{a^{\prime}\in\mathcal{A}}Q(s^{\prime},a^{\prime})]$.
@@
Preference-based reinforcement learning. Different from the standard RL setting, the reward function is not available in PbRL. Instead, PbRL learns the reward function $\widehat{R}$ from preferences between pairs of trajectory segments to align the human intention [15], where a trajectory segment $\sigma$ of length $L$ is defined as a sequence of states and actions $\{s_{k},a_{k},...,s_{k+L-1},a_{k+L-1}\}\in(\mathcal{S}\times\mathcal{A})^{L}$. Given a pair of segments $(\sigma_{0},\sigma_{1})$, human choose which segment is preferred, i.e., $y\in\{0,1\}$ [14]. The preference label $y=1$ indicates $\sigma_{1}\succ\sigma_{0}$ and $y=0$ indicates $\sigma_{0}\succ\sigma_{1}$, where $\sigma_{i}\succ\sigma_{j}$ denotes that the segment $i$ is preferable to the segment $j$.
@@
[…]
@@
In practical implantation, the transition model is approximated by Gaussian distribution and the MLE method is employed to train $N_{T}$ ensemble transition models $\{T^{i}_{\varphi}=\mathcal{N}(\mu^{i}_{\varphi},\sigma^{i}_{\varphi})\}_{i=1}^{N_{T}}$ [37]. The samples are generated through $H$-step rollouts. Here, we use the trained transition to generate more trajectories to achieve data augmentation.
III-B Problem Formulation
The cost of collecting preference data is high since it needs real-time human feedback. Therefore, this paper aims to improve sample efficiency to learn an accurate reward model from limited preference data and guarantee agent performance. For PbRL, the ideal form is that the learned reward function $\widehat{R}$ from collected preference data can be consistent with the true reward $R^{*}$. Here, we define the function class as $\mathcal{R}=\{R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}\}$. Then, similar to [44, 43], we make the below assumption for the reward model:
@@ III-B Problem Formulation
This assumption states that there must exist a reward function $R\in\mathcal{R}$ which can well-approximate the optimal reward function when the preference data can cover all data space. However, preference data tend to be limited, even fewer than a hundred in real-world scenarios. Therefore, different from previous PbRL [3, 35, 33], this paper aims to study how to learn the better reward function from the limited dataset $\mathcal{D}_{l}=\{(\sigma_{0}^{l},\sigma_{1}^{l},y)^{(i)}\}_{i=1}^{N_{l}}$. To broaden the coverage of preference data, motivated by model-based offline RL methods, we train the transition model through Eq. (4) to generate more unlabeled preference data $\mathcal{D}_{u}=\{(\sigma_{0}^{u},\sigma_{1}^{u})^{(i)}\}_{i=1}^{N_{u}}$. Then, the pseudo labels $\widehat{y}$ for unlabeled dataset $\mathcal{D}_{u}$ can be obtained through reward model trained on $\mathcal{D}_{l}$ [3], that is
@@
where $\mathbbm{1}(\cdot)$ is indicator function. The reward function can be updated through the collected labeled dataset $\mathcal{D}_{l}$ and the generated unlabeled dataset $\mathcal{D}_{u}$. Then, the reward model can be optimized by minimizing
@@
[…]
where $L((\sigma_{0},\sigma_{1}),y)=-(1-y)\log P(\sigma_{0}\succ\sigma_{1};{\psi})-y\log P(\sigma_{1}\succ\sigma_{0};{\psi})$. However, the pretrained reward model may generate incorrect pseudo-labels for the unlabeled dataset, leading to noisy training and poor generalization [31]. Therefore, one key question is how to design a data selection mechanism to improve prediction accuracy and guarantee training stability.
Another key aspect is the theory for offline PbRL. There are very few algorithms specifically designed for offline PbRL with strong guarantees, including a generalization bound for the reward model and a safe improvement guarantee for policy learning. Although Zhan et al.[18] established the lower bounds of policy improvement under partial coverage for offline PbRL, the general reward functions are defined over the whole trajectory rather than just state-action pairs, and the generalization bound of the reward model is not considered. Therefore, another key question is to develop a systematic theory for offline PbRL based on state-action pairs, including a generalization bound for the reward model and an improvement guarantee for policy learning.
IV Reward Learning
The reward learning involves two stages: pretraining based on collected labeled data $\mathcal{D}_{l}$ and updating based on $\mathcal{D}_{l}$ and unlabeled data $\mathcal{D}_{u}$ during policy learning, which is essentially semi-supervised learning [26]. This section focuses on designing a data selection function $f(\sigma_{0}^{u},\sigma_{1}^{u})$ to ensure the quality of generated preference data, and explaining the factors that influence the generalization ability of the reward model.
IV-A Uncertainty-aware Pseudo-label Selection
[…]